Thr angle between p vector and the resultant of p vector+ q vector and p vector -q vector

Grade:12th pass

1 Answers

Khimraj
5 years ago
resultant of (p+q) = (p-q) = 2p
so angle of p with 2p is zero.
Hope it clears..........................